### v1

# Shared DIFF Transformer

Abstract
DIFF Transformer improves attention allocation by enhancing focus on relevant context while suppressing noise. It introduces a differential attention mechanism that calculates the difference between two independently generated attention distributions, effectively reducing noise and promoting sparse attention patterns. However, the independent signal generation in DIFF Transformer results in parameter redundancy and suboptimal utilization of information. In this work, we propose Shared DIFF Transformer, which draws on the idea of a differential amplifier by introducing a shared base matrix to model global patterns and incorporating low-rank updates to enhance task-specific flexibility. This design significantly reduces parameter redundancy, improves efficiency, and retains strong noise suppression capabilities. Experimental results show that, compared to DIFF Transformer, our method achieves better performance in tasks such as long-sequence modeling, key information retrieval, and in-context learning. Our work provides a novel and efficient approach to optimizing differential attention mechanisms and advancing robust Transformer architectures.
Natural Language Processing, Transformers, ICML
1 Introduction
Transformers have achieved remarkable success across various tasks, from natural language processing to vision applications, largely due to their powerful self-attention mechanism. However, standard Transformers often overallocate attention to irrelevant context, leading to inefficiencies in both computational cost and model performance. This tendency becomes particularly pronounced in tasks requiring long-context modeling or precise key information retrieval, where irrelevant context can dominate attention distributions, hindering the model’s ability to focus on critical inputs.
To address this limitation, DIFF Transformer draws on the idea of noise-canceling headphones by introducing a differential attention mechanism that calculates the difference between two independently generated attention distributions. By amplifying relevant context and canceling out noise, DIFF Transformer improves attention allocation and promotes sparse attention patterns. This approach allows DIFF Transformer to focus on key inputs in long-context modeling and reduce interference from irrelevant context. However, the independent generation of attention signals in DIFF Transformer leads to parameter redundancy and suboptimal utilization of shared global information. While it effectively addresses noise suppression, there remains room for optimization in model complexity and computational efficiency.
To address this limitation, in this study, we propose the Shared DIFF Transformer, which draws on the idea of a differential amplifier by introducing a shared base matrix to model global patterns and incorporating low-rank updates to enhance task-specific flexibility. Similar to a differential amplifier, which calculates the difference between two signals to amplify the relevant signal and cancel out common-mode noise, Shared DIFF Transformer captures consistent global features through the shared base matrix, reducing parameter redundancy. Meanwhile, low-rank updates dynamically refine the two query matrices, amplifying meaningful signals while suppressing irrelevant noise. This design not only improves computational efficiency but also strengthens the robustness of differential attention in handling complex input scenarios.
We conducted extensive experiments across various tasks, including language modeling evaluation, scalability testing, long-context evaluation, key information retrieval, and in-context learning. The results show that Shared DIFF Transformer achieves comparable language modeling performance to DIFF Transformer while significantly reducing both the number of parameters and training tokens. In multiple downstream tasks, Shared DIFF Transformer not only demonstrates significant performance advantages but also exhibits excellent scalability. These findings position Shared DIFF Transformer as a robust and efficient architecture for large-scale language models, highlighting its effectiveness across various applications.
2 Shared Differential Transformer
2.1 Overview of Shared Differential Transformer
We propose the Shared Differential Transformer as a foundational architecture for sequence modeling, including large language models (LLMs). The architecture follows a decoder-only design and consists of $L$ Shared Differential Transformer layers. Given an input sequence $x=\{x_{1},x_{2},\dots,x_{N}\}$, we pack the input embeddings into
$X_{0}=[x_{1};x_{2};\dots;x_{N}]\in\mathbb{R}^{N\times d_{\text{model}}},$
where $d_{\text{model}}$ represents the hidden dimension of the model. Each layer contextualizes the input embeddings through the following operation:
$X_{l}=\text{Decoder}(X_{l-1}),\quad l\in[1,L],$
where each layer consists of two modules: a shared differential attention module and a feed-forward network (FFN) module. Additionally, we adopt pre-RMSNorm and SwiGLU, following best practices from models such as LLaMA, to improve stability and expressiveness.A diagram of the model architecture is shown in Figure 1.
Figure 1: Multi-head Shared DIFF Attention architecture.
(a) Scaling model size from 830M to 13B.
(b) Scaling the number of training tokens for the 3B model.
Figure 2: Language modeling loss when scaling model size and training tokens. Shared DIFF Transformer demonstrates superior performance, requiring fewer parameters or tokens to achieve similar results. (a) Shared DIFF Transformer achieves comparable performance to larger models while using fewer parameters. (b) Shared DIFF Transformer reaches similar performance with significantly fewer training tokens.
2.2 Shared Differential Attention
The shared differential attention mechanism maps query, key, and value vectors to outputs, leveraging a shared base matrix to model global patterns and low-rank updates for task-specific refinements. Specifically, given an input $X\in\mathbb{R}^{N\times d_{\text{model}}}$, we first project it to query, key, and value matrices as follows:
$Q_{1}=XW_{Q1},\quad Q_{2}=XW_{Q2},$ (1)
$K_{1}=XW_{K1},\quad K_{2}=XW_{K2},\quad V=XW_{V},$ (2)
where $W_{Q1},W_{Q2},W_{K1},W_{K2}\in\mathbb{R}^{d_{\text{model}}\times d}$ and $W_{V}\in\mathbb{R}^{d_{\text{model}}\times 2d}$.
To reduce parameter redundancy and enhance model flexibility, $W_{Q1},W_{Q2},W_{K1},W_{K2}$ are redefined using a shared base matrix and low-rank updates:
${\scriptsize W_{Q1}=W_{Q}+W_{q11}W_{q12}^{\top},\quad W_{Q2}=W_{Q}+W_{q21}W_{q22}^{\top}}$ (3)
${\scriptsize W_{K1}=W_{K}+W_{k11}W_{k12}^{\top},\quad W_{K2}=W_{K}+W_{k21}W_{k22}^{\top}}$ (4)
where $W_{Q}$ and $W_{K}\in\mathbb{R}^{d_{\text{model}}\times d}$ are shared base matrices used to capture global patterns. $W_{qi1}$ and $W_{ki1}\in\mathbb{R}^{d_{\text{model}}\times r}$, as well as $W_{qi2}$ and $W_{ki2}\in\mathbb{R}^{d\times r}$, are introduced as low-rank matrices for dynamic adjustments. These low-rank updates allow the model to adapt flexibly to different contexts while preserving shared global information, ensuring parameter efficiency and enhancing task-specific expressiveness.
The attention scores are computed as:
$A_{1}=\text{softmax}\left(\frac{Q_{1}K_{1}^{\top}}{\sqrt{d}}\right),\quad A_{2}=\text{softmax}\left(\frac{Q_{2}K_{2}^{\top}}{\sqrt{d}}\right).$ (5)
The final shared differential attention is defined as:
$\text{SharedDiffAttn}(X)=(A_{1}-\lambda A_{2})V,$
where $\lambda$ is a learnable scalar controlling the contribution of $A_{2}$. To stabilize learning dynamics, $\lambda$ is re-parameterized as:
$\lambda=\exp(\lambda_{q1}\cdot\lambda_{k1})-\exp(\lambda_{q2}\cdot\lambda_{k2})+\lambda_{\text{init}},$
where $\lambda_{q1},\lambda_{k1},\lambda_{q2},\lambda_{k2}\in\mathbb{R}^{d}$ are learnable vectors, and $\lambda_{\text{init}}\in(0,1)$ is a constant for initialization. This re-parameterization ensures consistent training dynamics, allowing us to effectively inherit hyperparameters from standard Transformers without extensive tuning.
2.3 Parameter Complexity
The introduction of shared base matrices significantly reduces parameter complexity compared to DIFF Transformer. Specifically, the parameter count for query and key projections is reduced from $4\cdot d_{\text{model}}\cdot d$ to $2\cdot d_{\text{model}}\cdot d+2\cdot d_{\text{model}}\cdot r+2\cdot d\cdot r$, where $r$ is the rank of the low-rank updates ($r\ll d,d_{\text{model}}$).
2.4 Multi-Head Shared Differential Attention
We extend the differential attention mechanism to support multi-head attention, which is a core feature of Transformer architectures. Let $h$ denote the number of attention heads, and $W_{i}^{Q}$, $W_{i}^{K}$, $W_{i}^{V}$, $i\in[1,h]$ be the projection matrices for each head. For the Shared Differential Transformer, the shared base matrices $W_{Q}$, $W_{K}$ are employed for all heads, with individual low-rank updates applied to each head to introduce head-specific variations.
The output of each attention head is computed as:
$\text{head}_{i}=\text{SharedDiffAttn}(X;W_{i}^{Q},W_{i}^{K},W_{i}^{V},\lambda),$
To ensure training stability, a scaling factor based on $\lambda_{\text{init}}$ is used to align the gradient flow with standard Transformers:
$\text{head}_{i}=(1-\lambda_{\text{init}})\cdot\text{LN}(\text{head}_{i}),$
where $\text{LN}(\cdot)$ denotes Layer Normalization. Aligning gradient flow ensures consistent training dynamics, enabling the effective use of hyperparameters from standard Transformer models. The final output of the multi-head shared differential attention is obtained by concatenating the outputs of all heads and projecting them:
$\text{MultiHead}(X)=\text{Concat}(\text{head}_{1},\dots,\text{head}_{h})W^{O},$
where $W^{O}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ is a learnable output projection matrix.
The number of heads is set to $h=\frac{d_{\text{model}}}{2d}$, where $d$ is the head dimension, ensuring alignment with the computational complexity and parameter count of standard Transformers.
2.5 Overall Architecture
The overall architecture of the Shared Differential Transformer consists of $L$ layers, where each layer includes a multi-head shared differential attention module followed by a feed-forward network (FFN) module. Given the input embeddings $X^{0}$, the $l$-th layer processes the output of the previous layer as:
$Y^{l}=\text{MultiHead}(\text{LN}(X^{l}))+X^{l},$
$X^{l+1}=\text{SwiGLU}(\text{LN}(Y^{l}))+Y^{l}.$
Here, $\text{LN}(\cdot)$ refers to RMSNorm, and $\text{SwiGLU}(\cdot)$ is defined as:
$\text{SwiGLU}(X)=\left(\text{swish}(XW^{G})\odot XW_{1}\right)W_{2},$
where $W^{G},W_{1}\in\mathbb{R}^{d_{\text{model}}\times\frac{8}{3}d_{\text{model}}}$, and $W_{2}\in\mathbb{R}^{\frac{8}{3}d_{\text{model}}\times d_{\text{model}}}$ are learnable matrices.
3 Experiments
We evaluate Shared DIFF Transformer from the following perspectives and compare it with DIFF Transformer. First, we compare the proposed architecture with DIFF Transformer in various downstream tasks and investigate the effects of scaling up model size and training tokens (Sections 3.1 and 3.2). Next, we present the results in key information retrieval and in-context learning, highlighting the advantages of Shared DIFF Transformer (Sections 3.3 and 3.4). Finally, we conduct extensive ablation studies to evaluate the impact of different design choices on model performance (Section 3.5).
3.1 Language Modeling Evaluation
Setup. We trained a 3B-size Shared DIFF Transformer language model and compared it with 3B-size DIFF Transformer language model. The model settings are shown in Table 1.
Params Values
Layers 30
Hidden size 2880
FFN size 7680
Vocab size 100,288
Heads 14
rank 256
Adam $\beta$ (0.9, 0.95)
LR $3.2\times 10^{-4}$
Batch size 4M
Warmup steps 1000
Weight decay 0.1
Dropout 0.0
Table 1: Configuration settings used for the 3B-size Shared DIFF Transformer models. Here, the ”rank” refers to the rank of the low-rank updates applied to the query and key matrices.
Results. Table 2 presents the results of zero-shot evaluation on the LM Eval Harness benchmark (Gao et al., 2021). We compare Shared DIFF Transformer with several state-of-the-art Transformer-based models, including OpenLLaMA-v2-3B (Geng & Liu, 2023), StableLM-base-alpha-3B-v2 (Tow, 2023), and StableLM-3B-4E1T (Tow et al., 2023). All models, including Shared DIFF Transformer, were trained on 1 trillion tokens under comparable conditions to ensure a fair comparison. The results clearly show that Shared DIFF Transformer outperforms these models in various downstream tasks, highlighting its enhanced ability to capture both local and global dependencies.
Notably, Shared DIFF Transformer consistently demonstrates improvements across multiple tasks, as reflected in the average score. The architectural enhancements, particularly the integration of the integral mechanism, allow for more effective utilization of global information. This significantly contributes to its strong performance on challenging tasks such as ARC-C, BoolQ, and PIQA.
Model ARC-C ARC-E BoolQ HellaSwag OBQA PIQA WinoGrande Avg
OpenLLaMA-3B-v2 33.9 67.6 65.7 70.0 26.6 76.7 62.9 57.5
StableLM-base-alpha-3B-v2 32.4 67.3 64.6 68.6 27.1 76.0 63.0 57.0
StableLM-3B-4E1T 66.6 25.5 76.8 63.2
DIFF-3B 37.8 72.9 69.0 71.4 29.0 76.8 67.1 60.6
Shared DIFF-3B 38.7 73.5 70.3 72.5 30.2 78.3 70.1 61.5
Table 2: Eval Harness accuracy compared with well-trained Transformer language models. The results indicate the superior performance of Shared DIFF Transformer over other models across a range of tasks.
3.2 Scalability Compared with Transformer
Size Hidden Dim. #Layers #Heads #Rank
830M 1536 24 8 96
1.4B 2048 24 10 128
2.8B 2880 32 14 256
6.8B 4096 35 20 320
13.1B 5120 44 30 448
Table 3: Model configurations for different sizes, including hidden dimension, number of layers, number of attention heads, and rank. Each model was trained with a sequence length of 2048 and a batch size of 0.25 million tokens, for a total of 40K training steps.
We assessed the scalability of Shared DIFF Transformer in comparison to the standard Transformer architecture, focusing specifically on language modeling tasks. This evaluation involved scaling both the model size and the number of training tokens. To ensure a fair comparison, we adopted a modified Transformer architecture similar to LLaMA, maintaining identical experimental setups across all models. The ”Transformer” models used in the comparison included optimizations like RMSNorm, SwiGLU, and the removal of biases.
Scaling Model Size As illustrated in Figure 2(a), Shared DIFF Transformer consistently outperformed both the Transformer and DIFF Transformer across a range of model sizes (for model configurations, refer to Table 3). Notably, Shared DIFF Transformer achieved similar validation loss to the Transformer while using 40% fewer parameters, and it matched the performance of DIFF Transformer with 24% fewer parameters. These results underscore the superior parameter efficiency and scalability of Shared DIFF Transformer.
Scaling Training Tokens Figure 2(b) presents the results from scaling the number of training tokens. The fitted curves show that Shared DIFF Transformer was able to achieve comparable performance to the Transformer with 30% fewer training tokens. Moreover, it surpassed the performance of DIFF Transformer with 11% fewer training tokens. These findings highlight the significant data efficiency of Shared DIFF Transformer, demonstrating its ability to deliver equivalent or superior performance with considerably fewer resources.
3.3 Key Information Retrieval
The Needle-In-A-Haystack test (Kamradt, 2023) is designed to evaluate how well models can identify key information in large contexts. In this test, ”needles” refer to brief sentences that link a city to a unique identifier. The goal is to accurately retrieve these identifiers from a query.
For the evaluation, we place the correct answer needle at various positions within the context (0%, 25%, 50%, 75%, 100%), while the remaining needles are placed randomly. Each combination of context length and needle position is evaluated over 50 samples, and the average accuracy is reported.
Model $N=1$ $N=2$ $N=4$ $N=6$
$R=1$ $R=2$ $R=2$ $R=2$
Transformer 1.00 0.85 0.62 0.55
DIFF 1.00 0.92 0.84 0.85
Shared DIFF 1.00 0.95 0.89 0.87
Table 4: Multi-needle retrieval accuracy in 4K-length contexts, averaged over the answer needle positions. $N$ represents the number of needles, and $R$ denotes the number of query cities.
Results from 4K Contexts We evaluated the multi-needle retrieval task using 4K-length contexts, with needle counts $N=1,2,4,6$ and retrieval counts $R=1,2$. All models were trained using 4K-length inputs. As shown in Table 4, Shared DIFF Transformer consistently outperformed both Transformer and DIFF models, particularly as the number of needles and query cities increased. For instance, with $N=6$ and $R=2$, Shared DIFF Transformer achieved an accuracy of 0.87, significantly surpassing the other models. This demonstrates that Shared DIFF Transformer excels in extracting relevant information even when surrounded by large amounts of irrelevant data, highlighting its robustness and efficiency in real-world tasks.
Retrieve from 64K Context Length As shown in Figure 3, we evaluated different context lengths with the configuration $N=8$, $R=1$. We assessed the 3B-size models with extended context (see Section 3.3). The results are reported across varying answer needle depths (y-axis) and context lengths (x-axis). From the results, it is evident that Shared DIFF Transformer outperforms both Transformer and DIFF Transformer. Notably, when the answer needle is placed at the 25% depth in a 40K context, Shared DIFF Transformer shows a 48% improvement over Transformer and an 8% improvement over DIFF Transformer in accuracy.
Figure 3: Multi-needle retrieval results in 64K length.
Attention Score Analysis Table 5 presents the attention scores assigned to the correct answer span and the irrelevant context in the key information retrieval task. These scores reflect how well the model focuses on relevant information while minimizing attention to noise. We compare the normalized attention scores at different depths (positions) of the target answer within the context. Compared to Transformer, Shared DIFF Transformer assigns significantly higher attention to the correct answer span and reduces attention to irrelevant context, especially in the early depths (0%, 25%, and 50%).
Model Attention to Answer↑ Attention Noise↓
0% 25% 50% 75% 100% 0% 25% 50% 75% 100%
Transformer 0.03 0.03 0.03 0.07 0.09 0.51 0.54 0.52 0.49 0.49
DIFF 0.27 0.30 0.31 0.32 0.40 0.01 0.02 0.02 0.02 0.01
Shared DIFF 0.33 0.36 0.39 0.41 0.44 0.01 0.01 0.02 0.01 0.01
Table 5: Attention scores allocated to answer spans and noise context in the key information retrieval task. The target answer is inserted at varying depths within the context. Shared DIFF Transformer allocates more attention to relevant information and effectively minimizes attention noise.
3.4 In-Context Learning
We explore in-context learning from two primary perspectives: its effectiveness in many-shot classification tasks and the model’s capacity to maintain robustness while leveraging context. In-context learning is a key feature of language models, highlighting their ability to efficiently utilize the given input context.
(a) TREC dataset (6 classes)
(b) TREC-fine dataset (50 classes)
(c) Banking-77 dataset (77 classes)
(d) Clinic-150 dataset (150 classes)
Figure 4: Accuracy of many-shot in-context learning across four datasets, with demonstration examples increasing from 1-shot up to a total of 64K tokens. The dashed lines indicate the average accuracy once the model’s performance stabilizes.
(a) Examples are randomly arranged.
(b) Examples are arranged alternately by class.
Figure 5: Many-shot in-context learning accuracy on four datasets. The accuracy for both DIFF Transformer and DINT (Ours) models is presented, showing performance improvements across different numbers of demonstration samples.
Model #Heads d GN Valid. Set↓ AR-Hit↓ Others↓
DIFF 8 256 3.062 0.880 3.247
−GroupNorm 8 128 3.122 0.911 3.309
with $\lambda_{\text{init}}=0.8$ 8 128 3.065 0.883 3.250
with $\lambda_{\text{init}}=0.5$ 8 128 3.066 0.882 3.251
Shared DIFF 8 128 3.057 0.876 3.245
−GroupNorm 8 128 3.110 0.903 3.297
with $\lambda_{\text{init}}=0.8$ 8 128 3.060 0.881 3.246
with $\lambda_{\text{init}}=0.5$ 8 128 3.059 0.881 3.247
Table 6: Evaluation of robustness in in-context learning on the TREC dataset.
Many-Shot In-Context Learning As shown in Figure 4, we compare the accuracy of DIFF Transformer and our Shared DIFF Transformer architecture in many-shot classification tasks. We evaluate the 3B-size language models with 64K input length support (see Section LABEL:sec:64k_length). We follow the evaluation protocol from (Bertsch et al., 2024) and use constrained decoding (Ratner et al., 2023). The number of demonstration samples is incrementally increased from 1-shot until the total length reaches 64K tokens. Specifically, we evaluate the models on the following datasets: TREC (Hovy et al., 2001) with 6 classes, TREC-Fine (Hovy et al., 2001) with 50 classes, Banking-77 (Casanueva et al., 2020) with 77 classes, and Clinic-150 (Larson et al., 2019) with 150 classes. The results show that Shared DIFF Transformer consistently outperforms DIFF Transformer across all datasets and varying numbers of demonstration samples. The improvement in average accuracy is significant, with Shared DIFF Transformer achieving 2.0% higher accuracy on TREC, 3.3% on TREC-Fine, 3.1% on Banking-77, and 1.5% on Clinic-150.
Robustness of In-Context Learning Figure 5 compares the robustness of DIFF Transformer and Shared DIFF Transformer in in-context learning. By analyzing how performance varies with order permutations of the same set of demonstration examples, we find that smaller performance fluctuations indicate greater robustness and a reduced risk of catastrophic degradation. The evaluation protocol follows the same methodology as previously described. Figure 5 shows the analysis results on the TREC dataset. We evaluate two prompt configurations: randomly shuffled examples and examples arranged alternately by class. In both configurations, Shared DIFF Transformer consistently exhibits smaller performance fluctuations compared to DIFF Transformer, demonstrating that our approach enhances robustness in in-context learning tasks. Specifically, in Figure 5.a, the fluctuation of Shared DIFF Transformer is reduced by 34%, and in Figure 5.b, the fluctuation is reduced by 42%.
3.5 Ablation Studies
We conduct ablation studies using the same training setup as described for the 1.4B model in Section 3.2. Table 6 reports the fine-grained loss on the validation set, breaking it into two components: ”AR-Hit” and ”Others.” ”AR-Hit” evaluates the model’s ability to recall previously seen n-grams, while ”Others” represents tokens that are either frequent or not recalled from the context.
As shown in Table 6, we conducted ablation studies on various design choices in DIFF Transformer and Shared DIFF Transformer and compared their performance. The first and fifth rows correspond to the default settings for DIFF Transformer and Shared DIFF Transformer, respectively. The results clearly show that GroupNorm has a significant impact on the model’s performance, as the rows in the attention matrices of both models no longer sum to 1. GroupNorm ensures numerical stability in this case.
We also experimented with different $\lambda$ initialization strategies. The results indicate that the default initialization method outperforms the constant initializations of $\lambda_{\text{init}}=0.8$ and 0.5. Moreover, the performance gap is minimal, highlighting the effectiveness of the initialization strategy and the robustness of the model to different initialization methods.
4 Conclusions
In this study, inspired by differential amplifiers, we propose Shared DIFF Transformer, which effectively reduces parameter complexity by introducing shared base matrices and low-rank updates. Through a series of extensive experiments, we validate the advantages of Shared DIFF Transformer across multiple natural language processing tasks, including text classification, question answering, and sequence generation. The experimental results demonstrate that Shared DIFF Transformer not only improves accuracy and recall rates but also exhibits stronger robustness and scalability. These advantages position Shared DIFF Transformer as a promising approach for future applications in natural language processing.